What is the formula for finding interest using natural logs and e? What is the difference between compounding continuously and compounding anually?
Answer by stanbon(57967)

What is the formula for finding interest using natural logs and e?
---
The formula for compounding continuously is A(t) = P*e^(rt)
-----------------------------------------------------------------
What is the difference between compounding continuously and compounding anually?
---
Annual compounding is A(t)=P(1+(r/n))^(nt) where n = 1
---
The Compounding continuously formula results from taking
the limit of (1+(r/n)^(nt) as n approaces infinity.
=======================================================
Cheers,
Stan H.
==============